OpenStudy (anonymous):

Use the squared identities to simplify 2sin^2 x cos^2 x. A. 1-cos(4x)/4 B. 1+4cos(2x)+cos(4x)/4 C. 1+cos(4x)/4 D. 1+4cos(2x)-cos(4x)/4

OpenStudy (raden):

hint : sin^2 x = (1-cos2x)/2 cos^2 x = (1+cos2x)/2

OpenStudy (raden):

well, i want to figure out this :